# Efficient Retrieval with Learned Similarities

### v2

@@ 4.2 Top $K$ retrieval performance
ML-1M BruteForce 1.00 1.00 1.00 1.00 1.00 .64$\pm$.03
TopKPerEmbd5 .838 .768 .681 .499 .437 .95$\pm$.03
TopKPerEmbd10 .932 .905 .842 .678 .618 .94$\pm$.05
TopKPerEmbd50 1.00 .997 .999 .977 .954 .95$\pm$.05
TopKPerEmbd100 1.00 1.00 1.00 .995 .990 1.24$\pm$.04
TopKAvg200 .980 .970 .967 .958 .948 .74$\pm$.06
TopKAvg500 .992 .986 .989 .990 .990 .71$\pm$.04
TopKAvg1000 .996 .994 .998 .997 .999 .72$\pm$.03
CombTopK5_200 1.00 .997 .988 .975 .960 1.09$\pm$.08
CombTopK50_500 1.00 .999 .999 .998 .997 1.13$\pm$.04
ML-20M BruteForce 1.00 1.00 1.00 1.00 1.00 3.08$\pm$.02
TopKPerEmbd5 .755 .702 .645 .467 .402 1.19$\pm$.05
TopKPerEmbd10 .946 .876 .817 .645 .564 1.17$\pm$.05
TopKPerEmbd50 .993 .986 .977 .931 .900 1.55$\pm$.03
TopKPerEmbd100 .998 .997 .994 .979 .966 2.26$\pm$.03
TopKAvg200 .992 .992 .992 .980 .958 .76$\pm$.04
TopKAvg500 .992 .993 .995 .995 .993 .77$\pm$.04
TopKAvg1000 .992 .994 .996 .998 .997 .75$\pm$.03
CombTopK5_200 1.00 .998 .998 .983 .962 1.32$\pm$.06
CombTopK50_500 1.00 1.00 1.00 .999 .996 1.85$\pm$.03
Books BruteForce 1.00 1.00 1.00 1.00 1.00 149.02$\pm$.49
TopKPerEmbd5 .907 .915 .809 .509 .396 20.17$\pm$.06
TopKPerEmbd10 .936 .986 .914 .698 .561 20.33$\pm$.06
TopKPerEmbd50 .993 .992 .994 .956 .902 21.09$\pm$.04
TopKPerEmbd100 1.00 .995 .996 .985 .959 22.53$\pm$.06
TopKAvg200 1.00 .978 .948 .845 .767 .68$\pm$.03
TopKAvg500 1.00 .992 .996 .919 .875 .68$\pm$.03
TopKAvg1000 1.00 1.00 .996 .963 .939 .84$\pm$.01
TopKAvg2000 1.00 1.00 1.00 .980 .968 1.04$\pm$.02
TopKAvg4000 1.00 1.00 1.00 .996 .987 1.64$\pm$.04
CombTopK5_200 .979 .984 .981 .892 .818 20.29$\pm$.05
CombTopK50_500 .993 .989 .994 .975 .961 21.57$\pm$.06
CombTopK100_1000 1.00 .997 .996 .993 .992 23.00$\pm$.08
Table 4: Evaluation of top $K$ retrieval performance, with hit rate (HR) normalized by the brute-force top $K$ method and latency with standard deviation measured over the evaluation of retrieval for a batch of queries (where the batch size is 32). (Relative) hit rate higher than .99 is marked in bold.
We evaluate the following methods on their top $K$ retrieval performance:
@@
Table 4 shows the hit rate (HR) and latency of all the methods. The hit rate is normalized by the ground truth, i.e., the hit rate achieved with brute-force top $K$. We measure latency by evaluating a batch of retrieval queries, in order to achieve high accelerator utilization; this is consistent with prior work on GPU/TPU-based retrieval algorithms [19, 4, 42]. We set batch size to $32$ for ML-1M, ML-20M, and Books. We perform evaluation on a single NVIDIA RTX 6000 Ada GPU. We report latency averaged over 20 warm runs.
@@
We observe that our approximate heuristics achieve high HR with oversampling. For example, TopKAvg500 is >.$>.99$ in relative HR across the board for ML-1M and ML-20M, and TopKAvg4000 is >.$>.99$ in relative HR across the board for Books. In addition, the combined top $K$ algorithm can outperform both TopKPerEmbd and TopKAvg of the corresponding configurations, sometimes significantly, e.g., CombTopK5_200 vs. TopKPerEmbd5 and TopKAvg200 on Books. This indicates that the set of candidate items retrieved by each individual approximate algorithm indeed complements each other when the weight distributions vary in MoL.
@@
In terms of efficiency, we observe that our approximate heuristics are significantly lower in latency than the exact baseline, especially as the number of items in the dataset becomes large. For example, compared to BruteForce, TopKAvg achieves >.$>.99$ relative HR@100 with a speedup of 4×4$\times$4 × and 91×91$\times$91 × in latency for ML-20M and Books, respectively. While the latency of the retrieval grows with the size of the dataset in the brute-force baseline, it grows much slower with the approximate methods. For example, the latency of retrieval increases by 4.8×4.8$\times$4.8 × from ML-1M to ML-20M in BruteForce, while the growth rate is 1.8×1.8$\times$1.8 × and $\times1$ for TopKPerEmbd100 and TopKAvg1000 respectively. Thus, we expect that the speedup of the approximate methods to become even more prominent as the size of the dataset further increases.
@@
We also notice that TopKAvg tends to be more efficient than TopKPerEmbd with comparable HR, e.g., TopKAvg2000 vs. TopKPerEmbd100 on Books with $\times2$ speedup in latency. We believe that this is mainly due to two reasons. First, when the HR is comparable, the maximal number of candidate items from TopKPerEmbd is larger than that of TopKAvg. Second, compared to TopKPerEmbd, the computation of TopKAvg is agnostic to the number of embedding sets because of the materialization optimization described in Section 3.2. Interestingly, we also see that the combined top $K$ is more efficient than the summation of the latency of its individual components, e.g., CombTopK5_200 is 1.5×1.5$\times$1.5 × faster than the sum of the latency from TopKPerEmbd5 and TopKAvg200 on ML-20M. This is because our implementation reduces the overhead of the combined method by consolidating some shared processing of the two components.
@@
Overall, empirically TopKAvg strikes a good balance between high HR and low latency, and the combined top $K$ can be used if the target HR is extremely stringent.
5 Related work
@@
Most information retrieval models in recommendations and natural language processing follow a classical two-stage paradigm [5], where up to billions of items [6, 26, 3] are first filtered down to hundreds in the retrieval stage, followed by the ranking stage that produces the final top 1-10 candidates. Earlier work on large-scale neural retrieval models primarily utilize dual-encoder (dense retrieval, etc.) setups, with dot products as the similarity function [5, 21]. Researchers quickly realized that dot products limited retrieval stage’s performance, and explored various learned similarity-based approaches. Prominent variants include maximum similarity based on multiple embeddings [26, 22, 34], specialized neural networks, often based on Hadamard products [14, 37, 39, 3], and representing item ids as token sequences (“learned index structures”), either implicitly defined during tree traversal [16, 45, 46] or explicitly in the “generative retrieval” setups [8, 38, 40, 36]. It has been shown, however, that learned neural distances often fail to outperform dot products, e.g., Hadamard MLPs in recommendations [33] and DSI in NLP [36]. Learned index structures further introduce stability and latency challenges as both NLP and recommendation systems need to support billion-scale realtime updated set of items [3]. Despite these challenges, significant gains (17% gains at Hit Rate@100 [42] to 24% gains at Hit Rate@400 [3]) with learned similarities have been reported in recent years; these can be attributed to careful construction of learned similarity functions [34, 42], implicit diversification done as part of beam search [8, 36], explicit representation of side-information using special neural architectures [3], and hardware-aware similarity function and inference algorithm design on GPUs [4, 42, 3].
Efficient Nearest Neighbor Search (NNS).
@@
Nearest neighbor search has been a popular topic of research due to their critical role in large-scale retrieval and vector databases. Most studies focus on the dot product case, also known as Maximum Inner Product Search (MIPS). Various techniques were proposed and analyzed, including tree structures [1, 32], locality sensitive hashing [10, 35], production quantization [18, 11], data partitioning [27, 44], graph-based methods [28, 17], and so on. The general case for NNS utilizing learned similarities remains less studied; for learned index structures, techniques to construct trees have been proposed to ensure beam search result in globally optimal top-$K$ results [46]. Algorithms based on implicit [28, 17, 39, 31] or explicit graphs [39] have been proposed to obtain a tractable candidate set in multi-stage retrieval setups; however, such approaches’ performance can degrade when the similarity function is not a metric, and constructing appropriate graph indices for non-metric similarity functions can remain challenging even for the inner product case [30]. Due to GPUs and other accelerators having orders of magnitude higher arithmetic intensity vs CPUs, traditional quantization techniques [35, 11] no longer fully utilize GPUs; accelerator-specific nearest neighbor algorithms that benefit from increased compute have been proposed and analyzed in recent work [4, 42, 31, 3].
6 Conclusion
@@
We have analyzed techniques for efficient retrieval with learned similarities in this work. We begin by showing Mixture-of-Logits (MoL) is a universal approximator of learned similarity functions, and further empirically learnable – MoL consistently outperforms dot products in HR across 6 scenarios by an average of 24.2% in HR@1 and 14.3% in HR@10, setting new state-of-the-art on recommendation datasets. We next propose both exact and approximate algorithms to enable efficient retrieval using learned similarity functions, and show their correctness and bounds. Across all datasets evaluated, we demonstrate that our approximate top $K$ algorithms can reach .99.99.99.99 of Hit Rate relative to exact algorithms, while achieving up to 91×91$\times$91 × reduction in end-to-end latency and with minimal indexing overheads. We expect the speedups to be further amplified with larger-scale datasets and GPU kernel optimizations. Given MoL’s empirical impressive performance gains of 20%-30% across Hit Rate@50-400 over hundreds of millions to billions of items [42, 3], our work concludes both theoretical and practical justifications for migrating vector databases away from dense retrieval and MIPS to Retrieval with Learned Similarities (RAILS) on GPUs.
Limitations
[…]
@@ Appendix A Experiment Setups
Our code is publicly available online, and detailed implementations and hyperparameter settings for reproducing our experiment results can be found at the following url: https://github.com/bailuding/rails. We discuss key settings below.
Parameterization of low-rank (“component-level”) embeddings, $f_{p}(q)$ and $g_{p}(x)$s for $1\leq p\leq P$.
@@
We note that embedding parameterization was not explicitly discussed in the original Mixture-of-Logits paper [42]. A specific implementation utilizing cluster id information has been proposed and discussed in a follow up work [3], where in the context of large-scale recommendation systems, the authors suggested cluster information based on interests of cohorts of members and topics of posts by themselves can lead to 10% recall gain at $K=400$. However, we cannot easily access similar information in the publicly available MovieLens and Amazon Reviews datasets. We therefore follow implementation provided by the authors [42] and additionally optionally utilizes a User ID keyed one-hot embedding as one query-side low-rank (“component-level”) embeddings $f_{p}(q)$, which is a widely used technique in recommendation systems [24]. All other $f_{p}(q)$s and $g_{p}(x)$s are obtained by applying an MLP on top of query-side/item-side representations in standard sequential recommendation setups [15, 20].
Parameterization of $\pi_{p}(q,x)$ matrices.

### v4

@@ 4.3. Top $K$ retrieval performance
ML-20M BruteForce 1.00 1.00 1.00 1.00 1.00 2.73$\pm$0.01
TopKPerEmbd5 0.69 0.67 0.63 0.46 0.40 1.61$\pm$0.05
TopKPerEmbd10 0.95 0.88 0.82 0.65 0.57 1.65$\pm$0.06
TopKPerEmbd50 1.00 1.00 0.99 0.95 0.92 1.64$\pm$0.05
TopKPerEmbd100 1.00 1.00 1.00 0.99 0.98 2.31$\pm$0.02
TopKAvg200 1.00 1.00 1.00 0.99 0.97 1.19$\pm$0.04
TopKAvg500 1.00 1.00 1.00 1.00 1.00 1.22$\pm$0.05
CombTopK5_200 1.00 1.00 1.00 1.00 1.00 1.86$\pm$0.06
Books BruteForce 1.00 1.00 1.00 1.00 1.00 128.36$\pm$0.30
TopKPerEmbd5 1.00 0.92 0.90 0.61 0.48 20.47$\pm$0.07
TopKPerEmbd50 1.00 1.00 1.00 0.98 0.93 21.60$\pm$0.08
TopKPerEmbd100 1.00 1.00 1.00 0.99 0.97 23.32$\pm$0.08
TopKPerEmbd200 1.00 1.00 1.00 1.00 1.00 26.55$\pm$0.12
TopKAvg200 0.97 0.92 0.91 0.76 0.67 1.13$\pm$0.04
TopKAvg500 0.97 0.98 0.97 0.86 0.81 1.17$\pm$0.04
TopKAvg1000 0.99 0.98 1.00 0.92 0.88 1.12$\pm$0.05
TopKAvg2000 1.00 0.99 1.00 0.95 0.92 1.20$\pm$0.02
TopKAvg4000 1.00 1.00 1.00 0.96 0.95 2.05$\pm$0.01
TopKAvg8000 1.00 1.00 1.00 0.97 0.97 3.79$\pm$0.01
CombTopK5_200 1.00 1.00 1.00 0.96 0.95 20.75$\pm$0.07
CombTopK50_500 1.00 1.00 1.00 0.99 0.96 22.12$\pm$0.07
CombTopK100_1000 1.00 1.00 1.00 0.99 0.98 24.02$\pm$0.13
CombTopK200_2000 1.00 1.00 1.00 1.00 1.00 28.01$\pm$0.11
NQ320K BruteForce 1.00 1.00 1.00 1.00 1.00 37.74$\pm$.47
TopKPerEmbd5 1.00 1.00 1.00 0.96 1.00 4.71$\pm$0.08
TopKPerEmbd10 1.00 1.00 1.00 0.98 1.00 4.83$\pm$0.08
TopKPerEmbd50 1.00 1.00 1.00 1.00 1.00 6.31$\pm$0.09
TopKAvg100 1.00 1.00 1.00 1.00 1.00 0.57$\pm$0.05
CombTopK5_100 1.00 1.00 1.00 1.00 1.00 5.28$\pm$0.08
Table 5. Evaluation of top $K$ retrieval performance, with hit rate (HR) normalized by the brute-force top $K$ method and latency measured over a batch of queries (where the batch size is 32). (Relative) hit rate higher than .99 is marked in bold.
We evaluate the following methods for top $K$ retrieval performance:
@@
For each dataset, we evaluate top $K$ retrieval methods based on the best performing model configurations reported in Table 3 and Table 4. Table 5 shows the hit rate (HR) and latency of all the methods. The hit rate is normalized by the ground truth, i.e., the hit rate achieved with brute-force top $K$. We measure latency by evaluating a batch of 32 retrieval queries, in order to achieve high accelerator utilization; this is consistent with prior work on GPU/TPU-based retrieval algorithms (Johnson et al., 2021; Chern et al., 2022; Zhai et al., 2023). We omit ML-1M as its size is small (Table 2). We set batch size to $32$ for all datasets. We perform evaluation on a single RTX 6000 Ada GPU. We report latency averaged over 20 warm runs.
@@
We observe that our approximate heuristics achieve high HR with overfetching. For example, TopKAvg500 achieves >.$>.99$ in relative HR across the board for ML-20M, and TopKAvg100 achieves >.$>.99$ in relative HR across the board for NQ320K. In addition, the combined top $K$ algorithm can outperform both TopKPerEmbd and TopKAvg of the corresponding configurations, sometimes significantly, e.g., CombTopK5_200 vs. TopKPerEmbd5 and TopKAvg200 on Books. This indicates that the set of candidate items retrieved by each individual approximate algorithm indeed complements each other when the weight distributions, $\pi_{p}(q,x)$s, vary in MoL.
@@
In terms of efficiency, we observe that our approximate heuristics are significantly lower in latency than the exact baselines, especially as the number of items in the dataset becomes large. For example, compared to BruteForce, TopKAvg achieves >.$>.99$ relative HR@100 with a speedup of $\times6$ in latency for NQ320K. While the algorithm latency grows with the size of the dataset in the brute-force baseline, it grows much slower with the approximate methods. For example, the algorithm latency increases by 47.0×47.0$\times$47.0 × from ML-20M to Books in BruteForce, while the growth rate is 10.1×10.1$\times$10.1 × and 1.0×1.0$\times$1.0 × for TopKPerEmbd100 and TopKAvg500, respectively. Thus, we expect the speedup of the approximate methods to become even more pronounced with larger datasets.
@@
We also notice that TopKAvg tends to be more efficient than TopKPerEmbd at comparable HRs, e.g., TopKAvg4000 is 10.5×10.5$\times$10.5 × faster than TopKPerEmbd50 on Books in terms of latency. First, the computation in TopKAvg is agnostic to the number of component-level embeddings, $P$, because of the preprocessing described in Section 3.2. Second, the branching and deduplication steps in TopKPerEmbd cannot leverage the parallel processing capabilities of GPUs effectively. Additionally, the latency of the combined top $K$ algorithm is lower than the sum of its individual components’ latencies; e.g., on ML-20M, CombTopK5_200 has a latency 1.5×1.5$\times$1.5 × lower than the sum of the individual latencies of TopKPerEmbd5 and TopKAvg200. This is done by consolidating processing shared by the two components.
@@
Overall, empirically TopKAvg strikes a good balance between high HR and low latency, and the combined top $K$ algorithm can be used if the target HR is extremely stringent.
5. Related work
@@
Most information retrieval models in recommendation systems and natural language processing (e.g., question answering) follow a classical two-stage paradigm (Covington et al., 2016; Karpukhin et al., 2020), where up to billions of items (Eksombatchai et al., 2018; Li et al., 2019; Zhai et al., 2023; Borisyuk et al., 2024) are first filtered down to hundreds in the retrieval stage, followed by another stage (e.g., ranking in recommendation systems or generation in RAG (Lewis et al., 2020)) that produces the final results. Earlier work on large-scale neural retrieval models primarily utilize dual-encoder (dense retrieval, etc.) setups, with dot products as the similarity function (Covington et al., 2016; Karpukhin et al., 2020; Ni et al., 2022a, b). Researchers quickly realized that dot products limited retrieval stage’s performance, and explored various learned similarity-based approaches. Prominent variants include maximum similarity based on multiple embeddings (Li et al., 2019; Khattab and Zaharia, 2020; Santhanam et al., 2022), specialized neural networks, often leveraging Hadamard products (He et al., 2017; Tan et al., 2020; Wang et al., 2022a; Borisyuk et al., 2024), and representing item ids as token sequences (“learned index structures”), either implicitly defined during tree traversal (Jasinska et al., 2016; Zhu et al., 2018; Zhuo et al., 2020) or explicitly in the “generative retrieval” setups (De Cao et al., 2021; Tay et al., 2022; Wang et al., 2022b; Bevilacqua et al., 2022; Zhuang et al., 2023; Sun et al., 2023). It has been shown, however, that learned neural distances often fail to outperform dot products, e.g., Hadamard MLPs in recommendation systems (Rendle et al., 2020) and DSI for QA scenarios in NLP (Sun et al., 2023). Learned index structures further introduce stability and latency challenges as both NLP and recommendation systems need to support billion-scale realtime updated set of items (Eksombatchai et al., 2018; Zhai et al., 2023; Borisyuk et al., 2024). Despite these challenges, significant gains (17% gains at Hit Rate@100 (Zhai et al., 2023) to 24% gains at Hit Rate@400 (Borisyuk et al., 2024)) with learned similarities have been reported in recent years; these can be attributed to careful construction of learned similarity functions (Santhanam et al., 2022; Zhai et al., 2023), implicit diversification done as part of beam search (Gao et al., 2021), explicit incorporation of side-information using special neural architectures (Zhai et al., 2023; Borisyuk et al., 2024), and hardware-aware similarity function and inference algorithm design on GPUs (Chern et al., 2022; Zhai et al., 2023; Ootomo et al., 2024; Borisyuk et al., 2024).
Load Balancing for Conditional Computations in Neural Networks
Conditional computations have been widely utilized in deep learning models (Bengio et al., 2016; Shazeer et al., 2017; Chen et al., 2023). Regularization losses have been proposed based on the observation that an ideal policy should evenly utilize all compute units in aggregate while being sparse at an individual example level (Bengio et al., 2016). Mixture-of-experts, a common way to implement conditional computations, has been widely used in language and vision domains (Chen et al., 2023; Shazeer et al., 2017) where mutual information-based regularization losses between experts and tasks (Chen et al., 2023) and experts and tokens (Shen et al., 2023) have been shown to help with various architectures.
Efficient Nearest Neighbor Search (NNS)
@@
NNS has been a popular research topic due to their critical role in large-scale retrieval and vector databases. Most studies focus on the dot product case, also known as Maximum Inner Product Search (MIPS). Various techniques were proposed and analyzed, including tree structures (Bentley, 1975; Ram and Gray, 2012), locality sensitive hashing (Gionis et al., 1999; Shrivastava and Li, 2014), production quantization (Jegou et al., 2011; Guo et al., 2016), data partitioning (Li et al., 2002; Zhai et al., 2011), graph-based methods (Malkov and Yashunin, 2020; Jayaram Subramanya et al., 2019), and so on. The general case for NNS utilizing learned similarities remains less studied; for learned index structures, techniques to construct trees have been proposed to ensure beam search result in globally optimal top-$K$ results (Zhuo et al., 2020). Algorithms based on implicit (Malkov and Yashunin, 2020; Jayaram Subramanya et al., 2019; Wang et al., 2022a; Ootomo et al., 2024) or explicit graphs (Wang et al., 2022a) have been proposed to obtain a tractable candidate set in multi-stage retrieval setups; however, such approaches’ performance can degrade when the similarity function is not a metric, and constructing appropriate graph indices for non-metric similarity functions can remain challenging even for the inner product case (Morozov and Babenko, 2018). Due to GPUs and other accelerators having orders of magnitude higher arithmetic intensity vs CPUs, traditional quantization techniques (Shrivastava and Li, 2014; Guo et al., 2016) no longer fully utilize the compute; accelerator-specific nearest neighbor algorithms that benefit from increased compute have been proposed recently (Chern et al., 2022; Zhai et al., 2023; Ootomo et al., 2024; Borisyuk et al., 2024).
6. Conclusion
@@
We have analyzed techniques for efficient retrieval with expressive learned similarities in this work. We begin by showing Mixture-of-Logits (MoL) is a universal approximator of all similarity functions, and further empirically learnable – MoL with our proposed load balancing loss consistently outperforms dot products (dense retrieval), sparse retrieval, and generative retrieval approaches across Recommendation Systems and Question Answering scenarios, setting new state-of-the-art across common, heterogeneous benchmark datasets. We next propose exact and approximate algorithms to enable efficient retrieval using learned similarity functions, and show their correctness and error bounds. Our approximate top $K$ algorithms can reach >.$>.99$ of Hit Rate relative to exact algorithms, while achieving up to $\times6$ reduction in end-to-end latency and with minimal indexing overheads. We expect the speedups to be further amplified with larger-scale datasets and GPU kernel optimizations. Given MoL’s impressive empirical performance gains of 20%-30% across Hit Rate@50-400 over hundreds of millions to billions of items (Zhai et al., 2023; Borisyuk et al., 2024) and broad applicability across heterogeneous scenarios, our work provides strong theoretical and practical justifications for migrating web-scale vector databases away from dense retrieval and MIPS to Retrieval with Learned Similarities (RAILS) on GPUs.
@@ A.1. Reproducibility
The implementations and hyperparameter settings for reproducing our experiment results can be found at https://github.com/bailuding/rails. We discuss specific details below.
A.2. Parameterization of low-rank (“component-level”) embeddings
In this section, we elaborate on the embedding parameterization methods for MoL that we discussed in Section 2.2.
A.2.1. Recommendation Systems
@@
Prior work have shown that careful parameterization of low-rank (“component-level”) embeddings, or $f_{p}(q)$ and $g_{p}(x)$s for $1\leq p\leq P$, can significantly improve MoL’s performance (Borisyuk et al., 2024). In the context of large-scale recommendation systems, cluster information based on interests of cohorts of members and topics of posts by themselves can lead to 10% recall gain at $K=400$ (Borisyuk et al., 2024). However, we cannot easily access similar information in the publicly available MovieLens (Harper and Konstan, 2015) and Amazon Reviews (McAuley et al., 2015) datasets. We therefore follow implementation provided by (Zhai et al., 2023) and additionally optionally utilizes a User ID keyed one-hot embedding as one query-side low-rank (“component-level”) embeddings $f_{p}(q)$, which is a widely used technique in recommendation systems (Koren et al., 2009) that we discussed in Section 2.2. All other component-level embeddings, $f_{p}(q)$s and $g_{p}(x)$s, are obtained by applying a multi-layer perceptron (MLP) on top of query-side/item-side representations in standard sequential recommendation setups (Hidasi et al., 2016; Kang and McAuley, 2018). The overall setup is illustrated on the right hand side of Figure 3.
Figure 3. Illustration of how to parameterize the embeddings to adapt Mixture-of-logits (MoL) learned similarity to various retrieval scenarios, with a language model (LM) finetuning use case in question answering (characterized by a single homogeneous feature) shown on the left, and a recommendation systems use case (characterized by a large number of heterogeneous features) shown on the right. For the Question Answering example on the left, $SP_{1},\ldots,SP_{N}$ represents the original SentencePiece (Kudo and Richardson, 2018) tokens that are inputs to the pre-trained language model LM, e.g., T5 (Raffel et al., 2023). $Q_{1},Q_{2},\ldots,Q_{P_{Q}}$ and $X_{1},X_{2},\ldots,X_{P_{X}}$ represent the special aggregation tokens we add to the LM tokenizer for pooling information across the sequence. The “Parameterized Pooling” component uses a $D$-dimensional embedding as input to parameterize, at an example-level, how to weight each of the (max_seq_len) encoder outputs for the $P_{Q}$/$P_{X}$ MoL component-level embeddings.
A.2.2. Question Answering (QA)
Unlike Recommendation Systems, retrieval models used in question answering generally take the full semantic representation(s) of the query and/or the document as input, and are finetuned on top of pre-trained language models with homogeneous inputs, or wordpiece / sentencepiece tokens. Our MoL embedding construction consists of two components, special aggregation tokens and parameterized pooling. We present embedding construction on the query side first.
Special Aggregation Tokens
Given both queries and documents are represented as token sequences (e.g., SentencePieces (Kudo and Richardson, 2018) in T5 (Raffel et al., 2023)), we propose to add special tokens that can be used to aggregate different aspects of information as part of the overall self-attention based language model. Specifically, on the query side, let the tokenized sequence be $SP_{1},SP_{2},\ldots,SP_{N}$. During finetuning of the pretrained language model, we create $P_{Q}$ special tokens, $Q_{1},\ldots,Q_{P_{Q}}$, and add them to the vocabulary of the query tokenizer. We also append those exact same $P_{Q}$ tokens before $SP_{1},SP_{2},\ldots,SP_{N}$, so that the $P_{Q}$ special tokens can be used to aggregate information across the query input using early-fusion mechanisms. Note that many question answering scenarios (Karpukhin et al., 2020; De Cao et al., 2021; Wang et al., 2022b; Ni et al., 2022b; Sun et al., 2023) utilize bidirectional language models for retrieval, like BERT (Devlin et al., 2019) or T5 (Raffel et al., 2023); for recent unidirectional language models, we can add the special aggregation tokens $X_{1},\ldots,X_{P_{X}}$ and $Q_{1},\ldots,Q_{P_{Q}}$ to the end of the input sequence instead. Our construction can also be viewed as a way to extend the CLS token in BERT (Devlin et al., 2019; Chang et al., 2023) to cover multiple aspects of information, in a way that encourages diversity via the $\mathcal{L}_{MI}$ load balancing loss discussed in Section 2.
Parameterized Pooling
We next add a pooling layer after the language model to encourage learning of aggregation mechanisms separate from language semantics. For each position $1\leq p\leq P_{Q}$, this pooling layer defines a probability distribution over different positions in language model’s outputs, or $(0,\ldots,max\_seq\_len-1)$. We further parameterize the pooling layer, using the $D$-dimensional embedding at the first position after encoders. This enables us to define a pooling policy, at an example-level, how to weight each of the $max\_seq\_len$ LM encoder outputs to arrive at the $P_{Q}$ MoL embeddings.
The embedding construction on the item-side is identical. We illustrate the overall finetuning setup we use for question answering on the left hand side of Figure 3.
A.3. Parameterization of $\pi_{p}(q,x)$ matrices